ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: Schmeelk, John, and Arpad Takaci, Virginia Commonwealth University Wavelet Transforms in Generalized Fock Spaces A generalized Fock space will be introduced as it was developed by Schmeelk and Takaci. The wavelet transform is then extended to this generalized Fock space. Since each component of our generalized Fock functional is a generalized function, the wavelet transform will act upon each individual entry much the same as was developed by Mikusinski and Mott. It is then shown that the generalized wavelet transform applied to a member of our generalized Fock space produces a more appropriate functional for certain applications.